General Chemistury 1420101 Quiz 1 Date: 20- 9 2elz Name: LD. Number:Lelkle43E Attendance Number etric relationships is incorrect? 1. Which of the follo 10 milliliters 1 liter C) 1 microliter 106 liters 10 decimeters 1 meter 1 gram 10 kilograms 2. The melting point ofa certain element is 343°C, What is this on the Fahrenheit scale? Ty Tex (9°F/5°C)+32°F) A) 890℉ B) 223°F C) 464 °F ) 649 °F The pressure of the earths atmosphere at sea level is 14.7 lb/in, what is the pressure when expressed in gm?? (2.54 cm = 1 in., 2.205 lb-1 kg) A) 4.30 x 10° g/m2 3· C) 5.02 x 1Og D) 1.03 x 10 g/m2 E) 2.09×10-5 g/m2 2.205 ㄛㄨㄧㄒ 4. An example of a pure substance is A) carbon dioxide B) compounds water E) all of these Page 1
Answer #1

Ans. #1. Incorrect relationships are – A, E

1 gram = 100 centigram

103 mL = 1 L

1 uL = 10-6 L

10 decimeter = 1 m

1 g = 10-3 kg = 0.001 kg

#2. Correct option E.

T(°F) = (T(°C) × 1.8) + 32 = 343 x 1.8 + 32 = 649.4 = 649

#3. Correct option D.

Given,

            2.205 lb = 1 kg

            So, 14.7 lb = (1 / 2.205) x 14.7 kg

= 6.666667 kg                                   ; [1 kg = 1000 g]

= 6666.667 g

            2.54 cm = 1 in

            Or, (1 in)2 = (2.54 cm)2 = 6.4516 cm2

                                    = 0.00064516 m2                            ; [1 cm2 = 0.0001 m2]

Now,

14.7 lb/ in2 = 6666.667 g / 0.00064516 m2

                        = 1.033 x 107 g/ m2

#4. Correct option. E. all of these

Any substance constituted of only one type of atom or molecules is called a pure substance.

Compounds, say CO2, H2O are pure substances because they consists of only one type of molecules, either CO2 or H2O, respectively.

Elements, say Na, or C, are pure substances because they consist of only one type of atoms.